# Data Reconstruction Attacks and Defenses: A Systematic Evaluation

### v1

@@ C.2 Differential Privacy
$\tilde{g}=g/\max\left\{1,\frac{\|g\|}{S}\right\}+\mathcal{E}_{g},$
where $\|G\|$ is the norm of the gradient of all parameters and $\mathcal{E}_{g}\sim\mathcal{N}(0,\sigma^{2}S^{2}I_{d})$, $d$ is the dimension of parameter. First, we give the proof of the case with no random noise that only gradient clipping is used, i.e. $\tilde{g}=g/\max{\left\{1,\frac{\|g\|}{S}\right\}}$. The error bound is the same as the case with no defense.
@@
For a 2-layer neural network with input dimension $d$, width $m$ and cross-entropy loss with $K$ classes, under Assumption A.1, we observe clipped gradient descent steps $\tilde{g}$ with batch size $B$, where $K^{2}B^{4}\leq\tilde{O}(d)$ and clipping threshold $S$. If $m\geq\tilde{\Omega}(\frac{K^{2}B^{2}d}{\nu^{2}\pi_{\min}^{2}})$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
For any $j$, we consider the last layer parameters for the first two classes $a_{1j}$ and $a_{2j}$. Let $g_{1j}=\sum_{i=1}^{B}\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{1j}}$ and $g_{2j}=\sum_{i=1}^{B}\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{2j}}$. Then the observed gradient updates are $\tilde{g}_{1j}=g_{1j}/\max\left\{1,\frac{\|G\|}{S}\right\}$ and $\tilde{g}_{2j}=g_{2j}/\max\left\{1,\frac{\|G\|}{S}\right\}$. Let $\tilde{g}({\bm{w}}_{j})=\tilde{g}_{1j}+\tilde{g}_{2j}$.
We define $\tilde{P}=\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. $\hat{P}$, $\hat{{\bm{T}}}$, $P$ and ${\bm{T}}$ are defined in the proof of Theorem B.1. Then let $R=\min\{1,\frac{S}{\|G\|}\}$, we have $\tilde{P}=R\hat{P}$ and $\tilde{{\bm{T}}}=R\hat{{\bm{T}}}$. Then by Eq. (10) and (12) we have
@@
with high probability by the proof of Theorem B.1.
On the other hand, the smallest components of $RP$ and $R{\bm{T}}$ are $R\nu_{\min}$ and $R\lambda_{\min}$ respectively. By the proof of Theorem B.1, $\nu_{\min}\geq\frac{|\nu|}{K}$, $\lambda_{\min}\geq\frac{|\lambda|}{K}$ and $\kappa\leq K$. Then by Theorem A.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}}_{i}\|\leq\tilde{O}(\frac{RB\sqrt{d/m}}{R|\mu|\pi_{\min}^{2}/K})+\tilde{O}(\frac{RK\sqrt{B^{6}/m}}{R|\lambda|\pi_{\min}^{2}/K})\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
Then we give the formal statement and proof of the error bound of the case without gradient clipping that the only defense is random noise, i.e. $\tilde{g}=g+\mathcal{E}_{g}$, where $\mathcal{E}_{g}\sim\mathcal{N}(0,\frac{\sigma^{2}}{m^{2}}I)$.
@@
For a 2-layer neural network with input dimension $d$, width $m$ and cross-entropy loss with $K$ classes, under Assumption A.1, we observe noisy gradient descent steps $g+\epsilon$ with batch size $B$, where $K^{2}B^{4}\leq\tilde{O}(d)$ and $\epsilon\sim\mathcal{N}(0,\frac{\sigma^{2}}{m^{2}}I)$. If $m\geq\tilde{\Omega}(\frac{K^{2}B^{2}d}{\nu^{2}\pi_{\min}^{2}})$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}((B+\sigma)\sqrt{\frac{d}{m}})$
@@
For any $j$, we consider the last layer parameters for the first two classes $a_{1j}$ and $a_{2j}$ and let $g_{1j}=\sum_{i=1}^{B}\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{1j}}$ and $g_{2j}=\sum_{i=1}^{B}\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{2j}}$. Then the observed gradient updates are $\tilde{g}_{1j}=g_{1j}+\frac{\epsilon_{1j}}{m}$ and $\tilde{g}_{2j}=g_{2j}+\frac{\epsilon_{2j}}{m}$, where $\epsilon_{1j},\epsilon_{2j}\sim\mathcal{N}(0,\sigma^{2})$. Let $\tilde{g}({\bm{w}}_{j})=\tilde{g}_{1j}+\tilde{g}_{2j}$.
We define $\tilde{P}=\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. $\hat{P}$, $\hat{{\bm{T}}}$, $P$ and ${\bm{T}}$ are defined in the proof of Theorem B.1. Then we have $\tilde{P}=\hat{P}+\frac{1}{m}\sum_{j=1}^{m}(\epsilon_{1j}+\epsilon_{2j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\hat{{\bm{T}}}+\frac{1}{m}\sum_{j=1}^{m}(\epsilon_{1j}+\epsilon_{2j})({\bm{w}}_{j}^{\otimes 3}+{\bm{w}}_{j}\tilde{\otimes}I)$. Now we bound $\|\tilde{P}-P\|$ and $\|\tilde{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|$.
Error bound of $\tilde{P}$. We have $\|\tilde{P}-P\|\leq\|\hat{P}-P\|+\|P_{\epsilon}\|,$ where $P_{\epsilon}=\frac{1}{m}\sum_{j=1}^{m}(\epsilon_{1j}+\epsilon_{2j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$. Since $\|\hat{P}-P\|\leq\tilde{O}(\frac{B\sqrt{d}}{\sqrt{m}})$ with high probability by Eq. (10), we only have to bound $\|P_{\epsilon}\|.$ We let $Z_{j}=(\epsilon_{1j}+\epsilon_{2j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and check the conditions of Theorem A.2:
@@
$\left\|Z_{j}\right\|\leq\left|\epsilon_{1j}+\epsilon_{2j}\right|\left\|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|\leq C\sigma d(\log(16m/\delta))^{2}$
@@
$\max\left\{\left\|\mathbb{E}[Z^{\top}Z]\right\|,\left\|\mathbb{E}[ZZ^{\top}]\right\|\right\}=\left\|\operatorname{\mathbb{E}}(\epsilon_{1}+\epsilon_{2})^{2}\operatorname{\mathbb{E}}[({\bm{w}}{\bm{w}}^{\top}-I)^{2}]\right\|.$
Let $Q=({\bm{w}}{\bm{w}}^{\top}-I)^{2}$. Since $Q_{ij}=\sum_{k\neq i,j}w_{i}w_{j}w_{k}^{2}+(w_{i}^{2}+w_{j}^{2}-2)w_{i}w_{j}$ for $i\neq j$ and $Q_{ii}=\sum_{k\neq i}w_{i}^{2}w_{k}^{2}+(w_{i}^{2}-1)^{2}$, $\operatorname{\mathbb{E}}({\bm{w}}{\bm{w}}^{\top}-I)^{2}=(d+1)I$. Then $\max\left\{\left\|\mathbb{E}[Z^{\top}Z]\right\|,\left\|\mathbb{E}[ZZ^{\top}]\right\|\right\}\leq O(\sigma^{2}d).$
@@
Then, $\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}\left(\operatorname{\mathbb{E}}\left[({\bm{a}}^{\top}Z\bm{b})^{2}\right]\right)^{1/2}=\leq O(\sigma^{2}).$
Moreover, $\operatorname{\mathbb{E}}[Z]=0.$ Then by Theorem A.2,
$\left\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}\right\|\leq\log(16m/\delta)\sqrt{\frac{\sigma^{2}d\log(4/\delta)}{m}}\leq\tilde{O}(\sigma\sqrt{\frac{d}{m}})$
@@
$\|\tilde{P}-P\|\leq\tilde{O}((B+\sigma)\sqrt{\frac{d}{m}})$
@@
Error bound of $\tilde{{\bm{T}}}(V,V,V)$. We have $\|\tilde{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|\leq\|\hat{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|+\|{\bm{T}}_{\epsilon}(V,V,V)\|,$ where ${\bm{T}}_{\epsilon}=\frac{1}{m}\sum_{j=1}^{m}(\epsilon_{1j}+\epsilon_{(}2j))({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. Since $\|\hat{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|\leq\tilde{O}(\frac{B^{5/2}}{\sqrt{m}})$ with high probability by Eq. (12), we only need to bound $\|{\bm{T}}_{\epsilon}(V,V,V)\|$. Note that $\|{\bm{T}}_{\epsilon}\|\leq\|T_{\epsilon}^{(1)}\|$, where $T_{\epsilon}^{(1)}$ is the flatten of ${\bm{T}}_{\epsilon}$ along the first dimension, so we can bound $\|T_{\epsilon}^{(1)}(V,V,V)\|$ instead. We check the conditions of Theorem A.2. For any ${\bm{x}}$
@@
$\left\|Z_{j}\right\|\lesssim|\epsilon_{1}+\epsilon_{2}|\|V^{\top}{\bm{w}}_{j}\|^{3}\leq\tilde{O}(\frac{\sigma B^{3/2}}{\sqrt{m}})$
@@
$\max\left\{\left\|\mathbb{E}\left[Z^{\top}Z\right]\right\|,\left\|\mathbb{E}\left[ZZ^{\top}\right]\right\|\right\}\leq\mathbb{E}\left[\left\|Z\right\|^{2}\right]\lesssim\mathbb{E}\left[(\epsilon_{1}+\epsilon_{2})^{2}\right]\mathbb{E}\left[\|V^{\top}{\bm{w}}_{j}\|^{6}\right]\lesssim\sigma^{2}B^{3}.$
@@
$\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}\left({\mathbb{E}}\left[\left({\bm{a}}^{\top}Z\bm{b}\right)^{2}\right]\right)^{1/2}\leq\left(\mathbb{E}\left[\left\|Z\right\|^{2}\right]\right)^{1/2}\lesssim\sigma B^{3/2}.$
Moreover, $\|\operatorname{\mathbb{E}}[Z]\|=0$. Then by Theorem A.2, we have for any $i$ that
$\|\frac{1}{m}\sum_{j=1}^{m}T^{(1)}_{\epsilon}(V,V,V)\|\leq\tilde{O}(\frac{\sigma B^{3/2}}{\sqrt{m}})$
@@
$\|\tilde{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|\leq\tilde{O}((B+\sigma)\sqrt{\frac{B^{3}}{m}})$
@@
Since $\nu_{\min}$, $\lambda_{\min}$ and $\kappa$ are not changed, by Theorem A.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}}_{i}\|\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}((B+\sigma)\sqrt{\frac{d}{m}})$
@@
Then we consider the case where both gradient clipping and gradient noise are used in the training. In this case, we found that gradient clipping will increase the error caused by gradient noise though clipping itself has no effect on the error bound. Here we reparameterize the observed gradient as $\tilde{g}=g/\max\left\{1,\frac{\|G\|}{S}\right\}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\frac{\sigma^{2}}{m^{2}}I)$.
Theorem C.12.
For a 2-layer neural network with input dimension $d$, width $m$ and cross-entropy loss with $K$ classes, under Assumption A.1, we observe clipped noisy gradient descent steps $g/\max\left\{1,\frac{\|G\|}{S}\right\}+\epsilon$ with batch size $B$, where $K^{2}B^{4}\leq\tilde{O}(d)$ and $\epsilon\sim\mathcal{N}(0,\frac{\sigma^{2}}{m^{2}}I)$. If $m\geq\tilde{\Omega}(\frac{K^{2}B^{2}d}{\nu^{2}\pi_{\min}^{2}})$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}((B+\sigma\max\{1,\frac{\|G\|}{S}\})\sqrt{\frac{d}{m}})$
@@
For any $j$, we consider the last layer parameters for the first two classes $a_{1j}$ and $a_{2j}$. Let $g_{1j}=\sum_{i=1}^{B}\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{1j}}$ and $g_{2j}=\sum_{i=1}^{B}\frac{\partial\ell(f({\bm{x}}_{i};\Theta),{\bm{y}}_{i})}{\partial a_{2j}}$. Then the observed gradient updates are $\tilde{g}_{1j}=g_{1j}/\max\left\{1,\frac{\|G\|}{S}\right\}+\epsilon_{ij}$ and $\tilde{g}_{2j}=g_{2j}/\max\left\{1,\frac{\|G\|}{S}\right\}+\epsilon_{2j}$, where $\epsilon_{1j},\epsilon_{2j}\sim\mathcal{N}(0,\sigma^{2})$. Let $\tilde{g}({\bm{w}}_{j})=\tilde{g}_{1j}+\tilde{g}_{2j}$.
We define $\tilde{P}=\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. $\hat{P}$, $\hat{{\bm{T}}}$, $P$ and ${\bm{T}}$ are defined in the proof of Theorem B.1. Let $R=\min\{1,\frac{S}{\|G\|}\}$. Then we have $\tilde{P}=R\hat{P}+\frac{1}{m}\sum_{j=1}^{m}(\epsilon_{1j}+\epsilon_{2j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=R\hat{{\bm{T}}}+\frac{1}{m}\sum_{j=1}^{m}(\epsilon_{1j}+\epsilon_{2j})({\bm{w}}_{j}^{\otimes 3}+{\bm{w}}_{j}\tilde{\otimes}I)$. By the proof of Proposition C.10 and Proposition C.11, we have
$\|\tilde{P}-RP\|\leq R\|\hat{P}-P\|+\|P_{\epsilon}\|\leq\tilde{O}((RB+\sigma)\frac{d}{m})$ (30)
@@
$\begin{split}\|\tilde{{\bm{T}}}(V,V,V)-R{\bm{T}}(V,V,V)\|&\leq R\|\hat{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|+\|{\bm{T}}_{\epsilon}(V,V,V)\|\\ &\leq\tilde{O}((RB+\sigma)\sqrt{\frac{B^{3}}{m}}).\end{split}$ (31)

### v2

@@ C.2 Differential Privacy
$\tilde{G}=G/\max\left\{1,\frac{\|G\|}{C}\right\}+\mathcal{E}_{G},$
where $\|G\|$ is the norm of the gradient of all parameters and $\mathcal{E}_{g}\sim\mathcal{N}(0,\sigma_{0}^{2}C^{2}I_{d})$, $d$ is the dimension of parameter. First, we give the proof of the case with no random noise that only gradient clipping is used, i.e. $\tilde{G}=G/\max{\left\{1,\frac{\|G\|}{C}\right\}}$. The error bound is the same as the case with no defense.
@@
Under Assumption B.1 and Assumption B.2, $y_{i}\in\{\pm 1\}$, we observe clipped gradient descent steps $\tilde{G}$ with batch size $B$, where $B^{4}\leq\tilde{O}(d)$ and clipping threshold $C$. If $m\geq\tilde{\Omega}(\frac{B^{2}d}{\nu^{2}\pi_{\min}^{2}})$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
For any $j$, the observed gradient updates for $a_{j}$ are $\tilde{g}({\bm{w}}_{j})=g({\bm{w}}_{j})/\max\left\{1,\frac{\|G\|}{C}\right\}$. We define $\tilde{P}=\frac{1}{m}\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\frac{1}{m}\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. $\hat{P}$, $\hat{{\bm{T}}}$, $P$ and ${\bm{T}}$ are defined in Eq. (6) to (9). Then let $R=\min\{1,\frac{C}{\|G\|}\}$, we have $\tilde{P}=R\hat{P}$ and $\tilde{{\bm{T}}}=R\hat{{\bm{T}}}$. Then by Lemma B.4 and B.5 we have
@@
with high probability.
On the other hand, the smallest components of $RP$ and $R{\bm{T}}$ are $R\nu_{\min}$ and $R\lambda_{\min}$ respectively. We have $\nu_{\min}\geq|\nu|$, $\lambda_{\min}\geq|\lambda|$ and $\kappa=1$. Then by Theorem B.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}}_{i}\|\leq\tilde{O}(\frac{RB\sqrt{d/m}}{R|\mu|\pi_{\min}^{2}})+\tilde{O}(\frac{R\sqrt{B^{6}/m}}{R|\lambda|\pi_{\min}^{2}})\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}(B\sqrt{\frac{d}{m}})$
@@
Then we give the formal statement and proof of the error bound of the case without gradient clipping that the only defense is random noise, i.e. $\tilde{G}=G+\mathcal{E}_{G}$, where $\mathcal{E}_{G}\sim\mathcal{N}(0,\sigma_{0}^{2}I)$.
@@
Under Assumption B.1 and Assumption B.2, $y_{i}\in\{\pm 1\}$, we observe noisy gradient descent steps $G+\epsilon_{0}$ with batch size $B$, where $K^{2}B^{4}\leq\tilde{O}(d)$ and $\epsilon\sim\mathcal{N}(0,\sigma_{0}^{2}I)$. If $m\geq\tilde{\Omega}(\frac{B^{2}d}{\nu^{2}\pi_{\min}^{2}})$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}((B+\sigma_{0})\sqrt{\frac{d}{m}})$
@@
For any $j$, the observed gradient updates for $a_{j}$ are $\tilde{g}({\bm{w}}_{j})=g({\bm{w}}_{j})+\epsilon_{j}$, where $\epsilon_{j}\sim\mathcal{N}(0,\sigma^{2})$. We define $\tilde{P}=\frac{1}{m}\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\frac{1}{m}\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. $\hat{P}$, $\hat{{\bm{T}}}$, $P$ and ${\bm{T}}$ are defined in Eq. (6) to (9). Then we have $\tilde{P}=\hat{P}+\frac{1}{m}\sum_{j=1}^{m}\epsilon_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\hat{{\bm{T}}}+\frac{1}{m}\sum_{j=1}^{m}\epsilon_{j}({\bm{w}}_{j}^{\otimes 3}+{\bm{w}}_{j}\tilde{\otimes}I)$. Now we bound $\|\tilde{P}-P\|$ and $\|\tilde{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|$.
Error bound of $\tilde{P}$. We have $\|\tilde{P}-P\|\leq\|\hat{P}-P\|+\|P_{\epsilon}\|,$ where $P_{\epsilon}=\frac{1}{m}\sum_{j=1}^{m}\epsilon_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$. Since $\|\hat{P}-P\|\leq\tilde{O}(\frac{B\sqrt{d}}{\sqrt{m}})$ with high probability by Lemma B.4, we only have to bound $\|P_{\epsilon}\|.$ We let $Z_{j}=(\epsilon_{1j}+\epsilon_{2j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and check the conditions of Theorem B.2:
@@
$\left\|Z_{j}\right\|\leq\left|\epsilon_{j}\right|\left\|({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)\right\|\lesssim\sigma_{0}d(\log(16m/\delta))^{2}$
@@
$\max\left\{\left\|\mathbb{E}[Z^{\top}Z]\right\|,\left\|\mathbb{E}[ZZ^{\top}]\right\|\right\}=\left\|\operatorname{\mathbb{E}}\epsilon^{2}\operatorname{\mathbb{E}}[({\bm{w}}{\bm{w}}^{\top}-I)^{2}]\right\|.$
Let $Q=({\bm{w}}{\bm{w}}^{\top}-I)^{2}$. Since $Q_{ij}=\sum_{k\neq i,j}w_{i}w_{j}w_{k}^{2}+(w_{i}^{2}+w_{j}^{2}-2)w_{i}w_{j}$ for $i\neq j$ and $Q_{ii}=\sum_{k\neq i}w_{i}^{2}w_{k}^{2}+(w_{i}^{2}-1)^{2}$, $\operatorname{\mathbb{E}}({\bm{w}}{\bm{w}}^{\top}-I)^{2}=(d+1)I$. Then $\max\left\{\left\|\mathbb{E}[Z^{\top}Z]\right\|,\left\|\mathbb{E}[ZZ^{\top}]\right\|\right\}\leq O(\sigma_{0}^{2}d).$
@@
Then, $\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}\left(\operatorname{\mathbb{E}}\left[({\bm{a}}^{\top}Z\bm{b})^{2}\right]\right)^{1/2}=\leq O(\sigma_{0}^{2}).$
Moreover, $\operatorname{\mathbb{E}}[Z]=0.$ Then by Theorem B.2,
$\left\|\frac{1}{m}\sum_{j=1}^{m}Z_{j}\right\|\leq\log(16m/\delta)\sqrt{\frac{\sigma_{0}^{2}d\log(4/\delta)}{m}}\leq\tilde{O}(\sigma_{0}\sqrt{\frac{d}{m}})$
@@
$\|\tilde{P}-P\|\leq\tilde{O}((B+\sigma_{0})\sqrt{\frac{d}{m}})$
@@
Error bound of $\tilde{{\bm{T}}}(V,V,V)$. We have $\|\tilde{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|\leq\|\hat{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|+\|{\bm{T}}_{\epsilon}(V,V,V)\|,$ where ${\bm{T}}_{\epsilon}=\frac{1}{m}\sum_{j=1}^{m}\epsilon_{j}({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. Since $\|\hat{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|\leq\tilde{O}(\frac{B^{5/2}}{\sqrt{m}})$ with high probability by Lemma B.5, we only need to bound $\|{\bm{T}}_{\epsilon}(V,V,V)\|$. Note that $\|{\bm{T}}_{\epsilon}\|\leq\|T_{\epsilon}^{(1)}\|$, where $T_{\epsilon}^{(1)}$ is the flatten of ${\bm{T}}_{\epsilon}$ along the first dimension, so we can bound $\|T_{\epsilon}^{(1)}(V,V,V)\|$ instead. We check the conditions of Theorem B.2. For any ${\bm{x}}$
@@
$\left\|Z_{j}\right\|\lesssim|\epsilon_{j}|\|V^{\top}{\bm{w}}_{j}\|^{3}\leq\tilde{O}(\frac{\sigma_{0}B^{3/2}}{\sqrt{m}})$
@@
$\max\left\{\left\|\mathbb{E}\left[Z^{\top}Z\right]\right\|,\left\|\mathbb{E}\left[ZZ^{\top}\right]\right\|\right\}\leq\mathbb{E}\left[\left\|Z\right\|^{2}\right]\lesssim\mathbb{E}\left[\epsilon^{2}\right]\mathbb{E}\left[\|V^{\top}{\bm{w}}_{j}\|^{6}\right]\lesssim\sigma_{0}^{2}B^{3}.$
@@
$\max_{\|{\bm{a}}\|=\|\bm{b}\|=1}\left({\mathbb{E}}\left[\left({\bm{a}}^{\top}Z\bm{b}\right)^{2}\right]\right)^{1/2}\leq\left(\mathbb{E}\left[\left\|Z\right\|^{2}\right]\right)^{1/2}\lesssim\sigma_{0}B^{3/2}.$
Moreover, $\|\operatorname{\mathbb{E}}[Z]\|=0$. Then by Theorem B.2, we have for any $i$ that
$\|\frac{1}{m}\sum_{j=1}^{m}{\bm{T}}^{(1)}_{\epsilon}(V,V,V)\|\leq\tilde{O}(\frac{\sigma_{0}B^{3/2}}{\sqrt{m}})$
@@
$\|\tilde{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|\leq\tilde{O}((B+\sigma_{0})\sqrt{\frac{B^{3}}{m}})$
@@
Since $\nu_{\min}$, $\lambda_{\min}$ and $\kappa$ are not changed, by Theorem B.1,
$\|{\bm{x}}_{i}-\hat{{\bm{x}}}_{i}\|\leq\frac{K}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}((B+\sigma_{0})\sqrt{\frac{d}{m}})$
@@
Then we consider the case where both gradient clipping and gradient noise are used in the training. In this case, we found that gradient clipping will increase the error caused by gradient noise though clipping itself has no effect on the error bound. Here we reparameterize the observed gradient as $\tilde{G}=G/\max\left\{1,\frac{\|G\|}{C}\right\}+\epsilon_{0}$, where $\epsilon_{0}\sim\mathcal{N}(0,\sigma^{2}I)$.
Proposition C.13 (Full version of Proposition 3.3).
Under Assumption B.1 and Assumption B.2, $y_{i}\in\{\pm 1\}$, we observe clipped noisy gradient descent steps $G/\max\left\{1,\frac{\|G\|}{C}\right\}+\epsilon$ with batch size $B$, where $B^{4}\leq\tilde{O}(d)$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$. If $m\geq\tilde{\Omega}(\frac{B^{2}d}{\nu^{2}\pi_{\min}^{2}})$, then with appropriate tensor decomposition methods and proper weights, we can reconstruct input data with the error bound
$\sqrt{\frac{1}{B}\sum_{i=1}^{B}\|{\bm{x}}_{i}-\hat{\bm{x}}_{i}\|^{2}}\leq\frac{1}{\min\{|\nu|,|\lambda|\}\pi_{\min}^{2}}\tilde{O}((B+\sigma_{0}\max\{1,\frac{\|G\|}{C}\})\sqrt{\frac{d}{m}})$
@@
For any $j$, the observed gradient updates for $a_{j}$ are $\tilde{g}({\bm{w}}_{j})=g({\bm{w}}_{j})/\max\left\{1,\frac{\|g\|}{C}\right\}+\epsilon_{j}$, where $\epsilon_{j}\sim\mathcal{N}(0,\sigma_{0}^{2})$. We define $\tilde{P}=\frac{1}{m}\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=\frac{1}{m}\sum_{j=1}^{m}\tilde{g}({\bm{w}}_{j})({\bm{w}}_{j}^{\otimes 3}-{\bm{w}}_{j}\tilde{\otimes}I)$. $\hat{P}$, $\hat{{\bm{T}}}$, $P$ and ${\bm{T}}$ are defined in Eq. (6) to (9). Let $R=\min\{1,\frac{C}{\|G\|}\}$. Then we have $\tilde{P}=R\hat{P}+\frac{1}{m}\sum_{j=1}^{m}\epsilon_{j}({\bm{w}}_{j}{\bm{w}}_{j}^{\top}-I)$ and $\tilde{{\bm{T}}}=R\hat{{\bm{T}}}+\frac{1}{m}\sum_{j=1}^{m}\epsilon_{j}({\bm{w}}_{j}^{\otimes 3}+{\bm{w}}_{j}\tilde{\otimes}I)$. By the proof of Proposition C.11 and Proposition C.12, we have
$\|\tilde{P}-RP\|\leq R\|\hat{P}-P\|+\|P_{\epsilon}\|\leq\tilde{O}((RB+\sigma_{0})\frac{d}{m})$ (25)
@@
$\begin{split}\|\tilde{{\bm{T}}}(V,V,V)-R{\bm{T}}(V,V,V)\|&\leq R\|\hat{{\bm{T}}}(V,V,V)-{\bm{T}}(V,V,V)\|+\|{\bm{T}}_{\epsilon}(V,V,V)\|\\ &\leq\tilde{O}((RB+\sigma_{0})\sqrt{\frac{B^{3}}{m}}).\end{split}$ (26)